# EVA: An Embodied World Model for Future Video Anticipation

### v1

@@ 4.5 Multi-stage Training Strategy
To train EVA, we employ a three-stage process. In the first stage, we train the Video-Language Model (VLM) and Video-Domain Model (VDM) separately, and using datasets like COCO (Chen et al., 2015) and a curated subset of CC3M-595K (Sharma et al., 2018) in VLM. The second stage involves aligning the VLM with an instruction tuning dataset, transforming it into an embodied QA bot using multimodal instruction data from sources such as MIMIC-IT (Li et al., 2023), LLaVA (Liu et al., 2023), and VideoChatGPT (Maaz et al., 2024), along with our EVA instruction tuning dataset (introduced in Section 5). In the final stage, we adapt the entire pipeline and train the adapter and VDM Ensemble-LoRA for each task, using the VLM’s output hidden embedding as a condition control signal. This stage addresses the diverse visual distribution among EVA-Instruction datasets(EVA-Instruct) by tuning the Ensemble-LoRA and adapter to maximize generation quality. More detailed information on these stages is provided in the appendix and experiment.
4.6 Self-Ask inference
Last, we propose self-ask inference steps to achieve the embodied prediction task. Given the question and visual observation, we first forward the inference of the VLM and obtain the hidden language state. The hidden state was later forwarded to the generation adapter, which was used as a condition to generate the prediction video. After each round of generation, the model follows the self-ask strategy (Press et al., 2022), which rethinks by asking the model itself, as shown in Figure 3, to check its prediction is a completed video. If the last frame does not satisfy the task completion requirement, the model will autonomously generate the extension of the video based on the previous prediction.
@@ 5 Experiments
In this section, we give a comprehensive experiment to evaluate the multimodal understanding and generation ability of EVA on four meta-tasks. First, we assess EVA’s VQA reasoning abilities on the Action-Description task in Section 5.1, highlighting the knowledge gaps in existing Vision-Language Models (VLMs) regarding embodied scenes. Next, we provide a detailed comparison of generation performance in Section 5.2, including the video score metrics from EVA-Score on the Finish-Thinking task. In Sections 5.3 and 5.4, we compare EVA against four baselines using EVA-Score to represent the capabilities when facing How-To and Next-Step tasks. Both qualitative and quantitative results are presented throughout, with additional qualitative analyses available in Appendix B. More video demonstrations can be found on the anonymous page at https://sites.google.com/view/eva-public.
Datasets for EVA-Bench. EVA-Bench includes a curated collection of 125 high-quality samples from our EVA-Instruct dataset. These samples encompass real-world robots, simulated robots, and egocentric human daily activities. The benchmark is categorized based on meta-tasks. details are included in Appendix C.
Implementation details. We set up two kinds of models, EVA-Generator and EVA. EVA-Generator uses Dynamicrafter as the backbone and fully fine-tunes it on the EVA-Instruct. EVA constructs an end-to-end pipeline with fine-tuned ChatUniVi as our VLM backbone and EVA-Generator as our VDM backbone, in the middle, we use a generation adapter to align the feature embedding among two backbones and only train the adapter with the EVA-Instruct.
5.1 Action-Description Video Question Answering
Metrics. First, we present the QA task results from our EVA-Bench. Comparing the BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), SPICE (Anderson et al., 2016), and CLIP (Radford et al., 2021) scores as multimodal measures. Furthermore, to have a better word analysis, we use GPT-4o (OpenAI, 2024) as an automatic evaluator to obtain a GPT-4o score. The detailed metric description is included in Appendix A.5.
Main Results. In Table 1, Qwen2-VL-7B (Yang et al., 2024) and GPT-4o significantly outperform other models in zero-shot inference. However, under GPT-4o’s evaluation, while Qwen2-VL-7B remains the best-performing open-source model with a score of 29.58, the gap between the LLaVA-NeXT Liu et al. (2024a) series and Qwen2-VL-7B has noticeably narrowed.
@@ 5.1 Action-Description Video Question Answering
ChatUniVi-base (Jin et al., 2024) 0.0969 0.0179 0.0640 0.1497 0.0427 0.0636 27.49 9.03
LLAVA-Next-Interleaved (Li et al., 2024b) 0.0725 0.0152 0.0741 0.1174 0.0843 0.0982 29.63 26.94
LLAVA-Next-Video (Zhang et al., 2024c) 0.0717 0.0249 0.0642 0.1062 0.1267 0.0961 30.36 25.56
LLAVA-OneVision (Li et al., 2024a) 0.0874 0.0306 0.0591 0.1118 0.2172 0.1043 27.97 22.35
Minicpmv2-6-7b (Hu et al., 2024) 0.0672 0.0164 0.0572 0.0913 0.0404 0.0456 28.88 17.63
qwen2-vl-7b (Yang et al., 2024) 0.2484 0.1643 0.1434 0.3255 0.8914 0.2839 28.98 29.58
GPT-4o (OpenAI, 2024) 0.2651 0.1058 0.1671 0.2902 0.7355 0.3015 22.96 33.19
ChatUniVi-loRA 0.3007 0.1855 0.1054 0.3268 0.8245 0.2213 24.89 31.94
ChatUniVi-Full-Parameter 0.4105 0.1544 0.1809 0.4416 1.9012 0.3414 25.36 38.46
EVA 0.5735 0.5012 0.3095 0.5873 4.0139 0.5506 24.98 62.63
Compared to the zero-shot models, the ChatUniVi (fine-tuned on a 50K subset of our EVA instruction tuning dataset) has better performance across many metrics. Our EVA model remains the top performer, achieving an impressive score of 62.63 under GPT-4o’s evaluation. This group of comparisons showcases that the mixture of data we use in EVA outperforms fine-tuning on pre-trained weights, which supports our multi-stage training method.
5.2 Finish-Thinking Video Generation
@@
Metrics We evaluate the task-condition generation quality of several models. We compare the video generation on Subject Consistency (SC), Background Consistency (BC), Motion Smoothness (MS), Dynamic Degree (DD) (Huang et al., 2024), Fréchet Video Distance (FVD), and especially Goal Completion Estimation(GCE) from task completion evaluation. The quantitative experiments are separated into three groups, as shown in Table 2. On the text domain, Finish-Thinking also includes VQA accuracy comparison on the output of Yes/No, which is not included in this table.
Main Results We fine-tune the full UNet of Dynamicrafter on the embodied video dataset named Dynamicrafter-Tune and compare it with the origin Dynamicrafter and EVA-Generator that use Ensamble-LoRA. The EVA-Generator improves significantly in GCE and FVD, with scores of 86.83 and 177.28, respectively. These improvements are due to our effective LoRA design, which successfully adapts the base model to different scenes separately.
Then, we evaluate the video reconstruction ability of the understanding and generation union model. The input video is first converted into a text description by the VLM and then recreated using the first frame and the description. We compare different pairs of VLM+VDM with EVA. Among these frameworks, EVA excels in SC(97.11), BC(96.01), and MS(99.31) and achieves the lowest FVD(184.81) score. This experiment demonstrates the quality of EVA and highlights the efficiency of our multi-stage training and adaption strategy in preserving the generation quality.
@@
Table 2: Finish-Thinking Video Generation Quality Comparison. Subject Consistency(SC), Background Consistency(BC), Motion Smoothness(MS), dynamic degree(DD), aesthetic quality(AQ), Goal Completion Estimation(GCE), Fréchet Video Distance(FVD)
@@
Dynamicrafter (Xing et al., 2023) Image+Text 87.25 91.91 96.72 63.33 43.57 80.96 362.56
Dynamicrafter-Tune Image+Text 83.49 89.70 97.87 64.28 36.70 80.72 235.52
EVA-Generator Image+Text 95.74 95.11 99.09 50.00 41.07 86.83 177.28
ChatUniVi+Dynamicrafter Video 87.10 90.82 96.97 70.00 44.38 80.80 314.11
qwen2-vl-7b+Dynamicrafter Video 87.13 91.39 96.29 73.33 44.19 81.96 307.33
ChatUniVi+EVA-Generator Video 96.54 95.26 99.19 43.33 40.47 88.48 189.61
qwen2-vl-7b+EVA-Generator Video 96.13 95.48 99.15 50.00 41.17 85.87 193,89
LLAVA-OneVision+EVA-Generator Video 96.64 95.54 99.17 50.00 41.11 88.74 192.83
EVA-2Stage Video 96.68 95.82 99.17 36.66 41.30 90.19 185.89
EVA Video 97.11 96.01 99.31 46.67 41.72 89.09 184.81
@@
      HOW-TO       LLAVA-OneVision+EVA-Generator       33.81       59.99       46.90
      qwen2-vl-7b+EVA-Generator       41.54       60.40       50.97
      EVA-2Stage       85.51       63.82       74.67
      EVA       85.51       69.93       77.72
      Next-Step       LLAVA-OneVision+EVA-Generator       16.75       51.75       34.25
      qwen2-vl-7b+EVA-Generator       42.99       57.23       50.11
      EVA-2Stage       73.02       62.10       67.56
      EVA       73.02       65.34       69.18
Figure 6: Interaction example of “use the kettle.” For each step, input the new observation video, and EVA can generate the instruction video, teaching the user how to use the kettle as an MR handbook.
5.3 How-To Instruction Generation
In Table 3, which demonstrates significant improvements, EVA-2Stage achieved an EVAS-L score of 85.51, an EVAS-V score of 63.82, and an overall EVA-Score of 74.67. Our proposed model, EVA, outperformed all other approaches with an EVAS-V score of 69.93 and an EVA-Score of 77.72. These results highlight that the EVA model, with its multi-stage pretraining and cross-attention alignment, significantly enhances both predictive accuracy and generalization in the How-To task.
Furthermore, by comparing these results, we observed that aligning the output of VLM models to the language instruction format of embodied scenes is essential. Additionally, adopting an end-to-end approach further enhances the quality of video generation, demonstrating the flexibility and effectiveness of our proposed method. For the qualitative results, we demonstrate how EVA can drive an image with different prompts under real and simulation robots in Appendix B
5.4 Embodied Next-Step Anticaption
@@
As shown in Table 3, for the quantitative result of the Next-Step task, EVA once again achieved the best performance. As a result, it provides a better text (semantic) condition to guide EVA-Generator for improved generation outcomes. In contrast, LLAVA-OneVision (Li et al., 2024a) and Qwen2-VL-7B performed worse in this task compared to the How-To scenario due to their inability to accurately predict the next-step description. This clearly demonstrates the importance of a VLM that is thoroughly trained in embodied scenes for the Embodied World Model. Such deficiency in EVAS-Language also affects the generation quality, leading to the lower performance of LLAVA-OneVision(51.75) and Qwen2-VL-7B(57.23) on EVAS-Vision. This comparison also shows the overall consistency and quality of EVA-Score.

### v2

@@ 4.2 Multi-stage Training Strategy
To train EVA, we employ a three-stage process. In the first stage, we use COCO (Chen et al., 2015) and a curate subset of CC3M-595K (Sharma et al., 2018) to warm up the visual encoder adapter in VLM. The second stage involves aligning the VLM with an instruction tuning dataset, transforming it into an embodied QA bot using multimodal instruction data from sources open-domain data e.g. VideoChatGPT (Maaz et al., 2024), along with our EVA instruction tuning dataset. In the final stage, we adapt the entire pipeline and train the generation adapter and VDM’s Ensemble-LoRA for each task. This stage addresses the diverse visual distribution among EVA-Instruction datasets by tuning the Ensemble-LoRA and adapter to maximize generation quality. We introduce detailed information on training stages is provided in the  Appendix A.
4.3 Autoregressive Frame Extention
The chunk-wise autoregressive frame extension method enables iterative video generation based on task completion verification. Given an input frame $o_{0}\in\mathbb{R}^{(B\times 1\times C\times H\times W)}$, the model first processes it through the VLM to obtain a hidden language state, which is passed to the generation adapter and used as a condition for the VDM to generate an initial video $v^{(0)}\in\mathbb{R}^{(B\times T\times C\times H\times W)}$. After each generation step, the model employs RoG by prefixed prompts as shown in  Figure 3 to verify whether the generated video satisfies the task’s completion requirement. If the task is incomplete, the model extracts the last 1 and $t$ frames ($v^{(k)}_{T-t,T-1}$) from the current output and uses them as input for the next round of generation, producing an extended video $v^{(k+1)}$. This process iterates, concatenating successive predictions ($v^{(0)},v^{(1)},\dots,v^{(k)}$) until the VLM determines that the task is complete. By iteratively refining predictions and focusing only on keyframes for extension, this method ensures coherent and task-aligned video generation.
5 Embodied Video Anticipation Benchmark
Under the RoG setting, we divide the complex interleaved multimodal generation task into four sub-tasks. This section describes how we separate the tasks and introduces our proposed dataset and benchmark. The fine-grained description of each metric and benchmark composition is provided in  Appendix C.
5.1 Dataset Construction
We construct four datasets for embodied tasks. The How-To and Action-Description datasets are built by extending prompts using GPT-4o (OpenAI, 2024) and standardizing annotations into a subject-verb-object format, based on the structure. For the Finish-Think dataset, we use the first 25% of videos to identify unfinished tasks and convert relevant RoboVQA (Sermanet et al., 2024) questions into this format. The Next-Step dataset is created using key step annotations from the Ego-Exo4D (Song et al., 2024) dataset and sequential task annotations from RoboVQA (Sermanet et al., 2024), focusing on ordered steps for next-step predictions. The complete EVA instruction tuning dataset comprises 500K QA pairs. Detailed information, including prompt structures, dataset ratios, and data examples, is provided in  Appendix C and anonymous supplementary pages.
5.2 Benchmark Construction
To ensure comprehensive evaluation, EVA-Bench includes 125 meticulously curated high-quality samples from the datasets introduced above, spanning diverse domains such as real-world robots, simulated robots, and egocentric human daily activities. These samples cover a wide range of scenarios, including pick-and-place tasks, cooking, bike repair, COVID testing, and indoor organization, providing a balanced distribution of tasks. For Finish-Thinking frames, we manually annotate the task completion frames. For out-of-distribution (OOD) evaluation in robotics, experts carefully annotate the prompts. Additionally, we rigorously annotate and refine all 125 samples and their corresponding prompts to ensure accuracy and consistency, making EVA-Bench a robust benchmark for embodied video anticipation.
@@ 6 Experiments
ChatUniVi (Jin et al., 2024) × 0.0969 0.0640 0.1497 0.0427 0.0636 27.49 9.03
LLAVA-NI (Li et al., 2024b) × 0.0725 0.0741 0.1174 0.0843 0.0982 29.63 26.94
LLAVA-NV (Zhang et al., 2024c) × 0.0717 0.0642 0.1062 0.1267 0.0961 30.36 25.56
LLAVA-O (Li et al., 2024a) × 0.0874 0.0591 0.1118 0.2172 0.1043 27.97 22.35
Minicpmv2 (Hu et al., 2024) × 0.0672 0.0572 0.0913 0.0404 0.0456 28.88 17.63
Qwen2 (Yang et al., 2024) × 0.2484 0.1434 0.3255 0.8914 0.2839 28.98 29.58
GPT-4o (OpenAI, 2024) × 0.2651 0.1671 0.2902 0.7355 0.3015 22.96 33.19
ChatUniVi-loRA 0.3007 0.1054 0.3268 0.8245 0.2213 24.89 31.94
ChatUniVi-FP 0.4105 0.1809 0.4416 1.9012 0.3414 25.36 38.46
EVA 0.5735 0.3095 0.5873 4.0139 0.5506 24.98 62.63
In this section, we give a comprehensive experiment to evaluate the multimodal understanding and generation ability of EVA on four meta-tasks. First, we assess EVA’s VQA reasoning abilities on the Action-Description task in Section 6.1, highlighting the knowledge gaps in existing VLMs regarding embodied scenes. Next, we provide the comparison of generation performance in Section 6.2. In Sections 6.3 we compare EVA against four baselines using EVA-Score to represent the capabilities when facing How-To and Next-Step tasks. Last, we demonstrate how the world model performs in robot tasks in  6.4. Both qualitative and quantitative results are presented throughout, with additional qualitative analyses available in Appendix B and page https://sites.google.com/view/icml-eva..
Datasets for Evaluation. EVA-Bench includes a curated collection of 125 high-quality samples from our EVA-Instruct dataset. These samples encompass real-world robots, simulated robots, and egocentric human daily activities. The benchmark is categorized based on meta-tasks. details are included in Appendix C.
Implementation details. We set up two kinds of models, EVA-Generator and EVA. EVA-Generator uses Dynamicrafter as the backbone and fully fine-tunes it on the EVA-Instruct. EVA constructs an end-to-end pipeline with fine-tuned ChatUniVi as our VLM backbone and EVA-Generator as our VDM backbone, in the middle, we use a generation adapter to align the feature embedding among two backbones and only train the adapter with the EVA-Instruct. We also implemented a discreet version of EVA-2stage by using the VLM and VDM in EVA without adapters.
6.1 Video Question Answering
EVA-Language Metrics. First, we present the QA task results from our EVA-Bench. Comparing the BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), SPICE (Anderson et al., 2016), and CLIP (Radford et al., 2021) scores as multimodal measures. Furthermore, to have a better word analysis, we use GPT-4o (OpenAI, 2024) as an automatic evaluator to obtain a GPT-4o score. The detailed metric description is included in Appendix A.7.
Main Results. In Table 1, we make the model fine-tuning on EVA-instruct as in-momain and compare a serial of zero-shot VLMs. Qwen2 (Yang et al., 2024) and GPT-4o excel in zero-shot inference. Under GPT-4o’s evaluation, while Qwen2 remains the best-performing open-source model with a score of 29.58, the gap between the LLaVA-NeXT (Liu et al., 2024a) series and Qwen2 has narrowed.
Fine-tuned models are trained on a 50K subset of our EVA instruction dataset by LoRA, or full-paramater(FP). Notably, our EVA model achieves the best performance with a score of 62.63 under GPT-4o’s evaluation, demonstrating that the diverse data and multi-stage training strategy used in EVA surpass fine-tuning on pre-trained weights.
6.2 Longer Video Generation
@@
EVA-Video Metrics. We use Subject Consistency (SC), Background Consistency (BC), Motion Smoothness (MS) (Huang et al., 2024), Fréchet Video Distance (FVD), and especially Goal Completion Estimation(GCE) for task completion evaluation. The quantitative experiments are separated into three groups, as shown in Table 2. Finish-Thinking also includes VQA accuracy comparison on the output of Yes/No, which is not included in the table.
Baselines For the model with only the video generation module, which takes image and text as input, we compare the Dynamicrafter (Xing et al., 2023), its fine-tuned version, and EVA’s video generation module(EVA-Gen). For the model with the understanding module under the RoG setting, we compared a series of 2-stage VLM+VDM models. We give the video as the input condition and ask the model to reconstruct. The input video is first converted into a text description by the VLM and then recreated using the first frame and the description.
Main Results For the video generation model with text and image as input, our fine-tuning version EVA-Gen improves significantly in GCE and FVD, with scores of 86.83 and 177.28, respectively, highlighting the effective LoRA design. Among these frameworks, EVA excels in SC(97.11), BC(96.01), and MS(99.31) and achieves the lowest FVD(184.81) score. This experiment demonstrates the quality of EVA and highlights the efficiency of RoG in task completion rate.
@@
Table 2: Finish-Thinking Video Generation Quality Comparison. Subject Consistency(SC), Background Consistency(BC), Motion Smoothness(MS), Goal Completion Estimation(GCE), Fréchet Video Distance(FVD). Blue means second-best results.
@@
DC (Xing et al., 2023) I+T 87.25 91.91 96.72 80.96 362.56
DC-Tune I+T 83.49 89.70 97.87 80.72 235.52
EVA-Gen I+T 95.74 95.11 99.09 86.83 177.28
ChatUniVi+DC V 87.10 90.82 96.97 80.80 314.11
Qwen2+DC V 87.13 91.39 96.29 81.96 307.33
ChatUniVi+EVA-Gen V 96.54 95.26 99.19 88.48 189.61
Qwen2+EVA-Gen V 96.13 95.48 99.15 85.87 193,89
LLAVA-O+EVA-Gen V 96.64 95.54 99.17 88.74 192.83
EVA-2Stage V 96.68 95.82 99.17 90.19 185.89
EVA V 97.11 96.01 99.31 89.09 184.81
@@
HOW-TO LLAVA-O+EVA-Gen 33.81 67.41 50.61
Qwen2+EVA-Gen 41.54 69.34 55.44
EVA-2Stage 85.51 73.32 79.42
EVA 85.51 77.83 81.67
Next-Step LLAVA-O+EVA-Gen 16.75 54.19 35.47
Qwen2+EVA-Gen 42.99 60.11 51.55
EVA-2Stage 73.02 64.46 68.74
EVA 73.02 68.68 70.85
6.3 Interleaved Generation
Among the two-stage methods, we demonstrate that EVA-L has a positive correlation with EVA-V. Understanding this ability can directly help improve generation quality and showcase the effectiveness of RoG. Our proposed model, EVA, outperformed other approaches under the RoG setting, achieving an EVAS-V score of 77.83 and an EVA-Score of 81.64. Qualitative results showcasing EVA’s performance with different prompts are included in Appendix B.
@@ 6.3 Interleaved Generation
As shown in Table 3, for the quantitative result of the Next-Step task, EVA had the best performance. As a result, it provides a better text (semantic) condition to guide EVA-Generator for improved generation outcomes. In contrast, LLAVA-OneVision (Li et al., 2024a) and Qwen2-VL-7B performed worse in this task compared to the How-To scenario due to their inability to accurately predict the next-step description. This clearly demonstrates the importance of a VLM that is thoroughly trained in embodied scenes for the Embodied World Model. Such deficiency in EVAS-L also affects the generation quality, leading to the lower performance of LLAVA-OneVision(51.75) and Qwen2-VL-7B(57.23) on EVAS-V. This comparison also shows the overall consistency and quality of EVA-S.
6.4 Evaluation on Robot Planning
Evaluation on RT1: We evaluated the success rates of human evaluation tasks in RT1 (Brohan et al., 2022) using the RoboDreamer (Zhou et al., 2024) framework, comparing AVDC, EVA without RoG, and EVA across in-domain and OOD prompts in  Table 4. EVA outperformed AVDC with a 28% higher overall success rate, achieving a 100% success rate in the Move Object task, demonstrating strong prompt-following ability. However, the Open/Close task proved particularly challenging due to cases like ”open right fridge door,” which involve transparent glass doors. In OOD, the EVA also shows an obvious improvement in successful counts, 5 more success cases than AVDC. The RoG contributes significantly, 7 cases higher than the model without RoG, demonstrating the importance of it’s frame extension.
Table 4: Task successful rate on in-domain and OOD tasks on RT-1. We randomly select six tasks from the RT1 validation set, spanning Pick Object, Move Object Near Object, Place Object Upright, Knock Object Over, Open/Close, Place Object into Receptacle.
Domain Model Pick Move to Place Knock Over Open/close Place into Sum.
In-Domain AVDC 11/16 9/12 1/2 4/4 2/8 2/8 29/50
w/o RoG 13/16 12/12 2/2 4/4 4/8 6/8 41/50
EVA 13/16 12/12 2/2 4/4 4/8 7/8 42/50
OOD AVDC 2/4 1/4 1/1 1/2 1/1 1/3 7/15
w/RoG 2/4 1/4 0/1 2/2 0/1 0/3 5/15
EVA 3/4 4/4 1/1 2/2 1/1 1/3 12/15
Evaluation on CALVIN We selected four tasks in CALVIN same as  (Luo & Du, 2024). As shown in  Table 5, EVA significantly outperforms the baseline (w/o RoG) in both video and simulation evaluations, demonstrating its superior robotic planning ability. On the video level, EVA achieves much higher task completion rates across all tasks, highlighting the effectiveness of RoG. In the simulation, we applied two versions of the video-to-action head, named CMLP and CMLP2. CMLP&CMLP2 uses video prediction and first action status as conditions, outputting 16 frames of actions. CMLP2 uses more tricks including simple action chunks, etc. The generated video The combination of EVA with CMLP2 achieves the best performance, significantly surpassing both the baseline and EVA with standard CMLP. These results confirm that both RoG and the enhanced CMLP2 model contribute to improved task planning and execution, and demonstrate the potential of this approach in robot task planning.
Table 5: Successful rate in CALVIN, including human evaluation of video, and task successful rate in simulation environments. The quality demos are included on the demo page.
eval model lightbulb led rotate Open drawer
Video w/o RoG 12/41 43/54 98/157 18/41
EVA 35/41 45/54 124/157 36/41
Simulation w/o RoG +CMLP 7/41 13/54 32/157 4/41
EVA+CMLP 9/41 15/54 45/157 9/41
w/o RoG+CMLP2 10/41 32/54 48/157 11/41
EVA+CMLP2 32/41 43/54 90/157 25/41